The ratio of boys to girls in a school is 5:2. The number of boys is more by
450 than that of girls. How many students are there in that school ?

Answer by TimothyLamb(4379)

---
x/y = 5/2
x = y + 450
---
put the system of linear equations into standard form:
---
2x - 5y = 0
x - y = 450
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = boys = 750
y = girls = 300
total students = 1050
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php